OpenStudy (anonymous):

is -2^2+3 a linear function? how can you tell?

OpenStudy (anonymous):

thats just a line

OpenStudy (lgbasallote):

i dont think it's a line...it doesnt even have a variable...

OpenStudy (anonymous):

a horizontal line if f(x)= what you typed. i think you forgot something

OpenStudy (lgbasallote):

ahh yes..now that's a line :D

OpenStudy (lgbasallote):

otherwise this isnt even a function

OpenStudy (anonymous):

just a constant

OpenStudy (lgbasallote):

just a number in the big black space

OpenStudy (anonymous):

the problem is "tell whether the function is linear. Evaluate when x=-2 ; F9x)=x^2+3

OpenStudy (anonymous):

i think they meant 2x^2+3 in that case you can prove for every x value there is a y value. making it a function of y in respect to x.

OpenStudy (lgbasallote):

x^2 + 3 is not linear though

OpenStudy (anonymous):

correct; it's parabolic

OpenStudy (anonymous):

so -2^2+3 it is linear then..?

OpenStudy (lgbasallote):

-2^2 + 3..for the nth time is NOT a function....we're talking about x^2 + 3 here...it is not linear...it is quadratic advise...in looking if a function is linear or not...do it before substituting...functions only exist if there is a variable

OpenStudy (freckles):

y=-2^2+3 is a constant so it is a line

OpenStudy (freckles):

x=-2^2+3 is vertical line but it not a function

OpenStudy (lgbasallote):

@freckles the question is "tell whether the function is linear. Evaluate when x=-2 ; F(x)=x^2+3" you tell whether the function is linear first THEN evaluate x = -2

OpenStudy (freckles):

For it to be a function it must pass the vertical line test and a vertical line doesn't do that